OpenStudy (anonymous):

.find the domain and range of the quadratic function and state whether the vertex is a maximum or minimum f(x)=-7(x-5)^2-7

OpenStudy (anonymous):

\[(x-5)^2\geq 0\] always, because it is a square that makes \[-7(x-2)^2\leq 0\] and so \[f(x)=9(x-5)^2-7\leq -7\]

OpenStudy (anonymous):

then do you graph is

OpenStudy (anonymous):

typo on the last line \[f(x)=-7(x-5)^2-7\leq -7\]

OpenStudy (anonymous):

vertex is \((5,-7)\) draw a parabola that opens down with that as the vertex

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

what about the domain and range